Circular No. 8049

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 A2 Arianna E. Gleason, Lunar and Planetary Laboratory, reports her dis-

covery of a slow-moving comet on Jan. 10.4 UT (discovery position below) with the Spacewatch II telescope at Kitt Peak; J. V. Scotti adds that there was a more-or-less symmetrical coma ∼ 20′′ across. On making followup observations on Jan. 11.3 (after placement on The NEO Confirmation Page), D. T. Durig and H. H. Fry (Sewanee, TN, 0.3-m f/5.75 Schmidt-Cassegrain reflector) confirmed a coma 15′′–18′′ in diameter, and F. B. Zoltowski (Edgewood, NM, 0.3-m f/3.3 Schmidt-Cassegrain reflector) noted that the coma/tail structure had the appearance of a broad fan from p.a. 20^{o} northward through p.a. 200^{o}. On Jan. 12.0, J. Tichá and M. Tichý (Kleť Observatory, 1.06-m KLENOT Telescope) indicated a coma diameter of 8′′–10′′, with m_{1} = 20.2 and m_{2} = 21.0. The object’s cometary nature was also noted by T. Gehrels (Spacewatch II) on Jan. 11–13 and by J. G. Ries (McDonald Observatory, 0.76-m reflector) on Jan. 14.3.

2003 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 10.39264  8^{h}55^{m}59.^{s}44  +13^{o}36′48′′.5  20.0

MPEC 2003-A78 contains the full astrometry and a very tentative parabolic orbit (assumed T = 2003 Jan. 10; hence q = 11.5 AU, i = 8^{o}). It seems likely that the object is a Centaur, showing cometary activity as (2060) = 95P/Chiron has shown near perihelion.

SUPERNOVA 2003I IN IC 2481 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003I (cf. IAUC 8046), obtained by M. Calkins on Jan. 12.41 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ib supernova, similar to SN 1998dt at 8 days past maximum (see Matheson et al. 2001, A.J. 121, 1648).

SUPERNOVA 2003H IN NGC 2207 Corrigendum. The final sentence on IAUC 8047 should be replaced

with: “The Na I D interstellar lines are present in absorption with an equivalent width of 0.6 nm, which suggests significant reddening in the host galaxy. Most likely this is the reason why the spectrum of SN 2003H is so red compared to that of the type-Ib supernova 1984L.”

2003 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green